OpenStudy (anonymous):

A fabric store has 80 3/8 yards of a particular fabric. About how many pairs of curtains could be made from this fabric if each pair requires 4 1/8 yards of fabric?

OpenStudy (compassionate):

Enough to never need fabric again. ;)

OpenStudy (australopithecus):

Total amount of fabric 80 + 3/8 yards 1 Curtain is 4 + 1/8 yards You can just use division to figure this out but it might help to express your total amount of fabric and the amount of fabric needed to make one curtain as a single fraction. Do you know how to do this?

OpenStudy (anonymous):

no I don't I'm that bad at math.. :(

OpenStudy (australopithecus):

To add two fractions both numbers must have the same denominator. In this case we have

OpenStudy (australopithecus):

\[\frac{80}{1} + \frac{3}{8}\]

OpenStudy (australopithecus):

Now we cannot add these numbers until they have the same denominator

OpenStudy (anonymous):

8? for the same denomiator

OpenStudy (australopithecus):

Yes, to do this we have to modify 80/1 to have a denominator of 8 without without changing its value

OpenStudy (australopithecus):

Remember the rule, that as long as you multiply the top and the bottom of a fraction by the same number it does not change its value. For instance \[\frac{3}{2} = \frac{3*2}{2*2} = \frac{\frac{3}{2}}{\frac{2}{2}}\]

OpenStudy (australopithecus):

Also remember that you can express any number with the denominator 1

OpenStudy (australopithecus):

so pretty much anything can be expressed as a fraction

OpenStudy (australopithecus):

The rule also goes with division as well as seen above

OpenStudy (australopithecus):

Do you follow?

OpenStudy (anonymous):

not really.. can you help me with this one? Estimate the difference 10 2/3-6 9/11 1) 5 2) 4 3) 3 4) 6 >.<

OpenStudy (australopithecus):

can you try to follow? This stuff is fundamental and I would rather teach you how to do this math than give you the answer

OpenStudy (anonymous):

I'm terrible at math I can't multiply or divide at all. :/ moved so much that it fluttered me up (sorry about the language..)

OpenStudy (australopithecus):

Fundamental you probably already know this but equal means that two things are the same thing for instance 5 = 5 five is the same thing as five 5 + 3 = 8 5 + 3 is the same as 8 Ok now moving onto your problem First off you need to understand and remember that in order to add or subtract fractions both fractions need to have the same denominator. Second you need to realize that any whole number can be expressed as a fraction (this will make it easier to write out), for instance \[3 = \frac{3}{1}\] \[\frac{5}{1} = 5\] \[555 = \frac{555}{1}\] This is because 1 goes into 5 five times, 1 goes into 3 three times. 1 goes into 555 five hundred and fifty five times etc. Secondly there is a rule that states, as long as you multiply or divide the numerator of a fraction by the same thing as you multiplied or divided the denominator of fraction, the fraction will remain equal to its original value. So, just say we wanted to give the number 6 the denominator of 5 Note: \[\frac{5}{5} = 1\] because 5 goes into 5 once Now onto our problem \[6 = \frac{6}{1}\] \[\frac{6}{1}*\frac{5}{5} = \frac{6*5}{1*5} = \frac{30}{5}\] \[\frac{30}{5} = 6\] which makes sense as 5 goes into 30 six times

OpenStudy (anonymous):

thank you...so very much <3 :)

OpenStudy (australopithecus):

Understanding how to manipulate fractions is vital to your success in mathematics please take this subject seriously and spend time learning it and mastering it

OpenStudy (australopithecus):

Once you have two fractions (numbers) with the same denominator you can add or subtract them for instance \[\frac{7}{4} + \frac{5}{4} = \frac{7 + 5}{4} = \frac{12}{4}\] it is the same with subtraction \[\frac{7}{4} - \frac{5}{4} = \frac{7 - 5}{4} = \frac{2}{4}\]

OpenStudy (australopithecus):

Whenever you see something like 10 3/4 or 4 2/3 it just means 10 + 3/4 and 4 + 2/3 once you have this written out you can combine them into one fraction using the rules I showed you above